Jessica walks and jogs to the university each day. If she averages 5 km/hr walking 9 km/hr jogging, how far does she jog in the distance from her home to the university is 8 km and Jessica makes the trip in 1 hour?
Answer by richwmiller(17219)

w+j=1,
5w+9j=8
j = 3/4, w = 1/4 hour
jog 9*3/4=27/4 km
walk 5*1/4=5/4 km
check
9*3/4+5/4=8
27/4+5/4=32/4
ok